OpenStudy (anonymous):

Triangle ABC has C=90º: Find angle B if a=4 and c=8.

OpenStudy (mathstudent55):



OpenStudy (anonymous):

use Pythagoras theorm b^c=c^2-a^2

OpenStudy (mathstudent55):

Since you are dealing with a right triangle, use the Pythagorean theorem.

OpenStudy (mathstudent55):

\(a^2 + b^2 = c^2\), so \(b^2 = c^2 - a^2\) Plug in values for a and c, and solve for b.